# Gotta Hear Them All: Sound Source-Aware Vision to Audio Generation

### v2

@@ Abstract
Vision-to-audio (V2A) synthesis has broad applications in multimedia. Recent advancements of V2A methods have made it possible to generate relevant audios from inputs of videos or still images. However, the immersiveness and expressiveness of the generation are limited. One possible problem is that existing methods solely rely on the global scene and overlook details of local sounding objects (i.e., sound sources). To address this issue, we propose a Sound Source-Aware V2A (SSV2A) generator. SSV2A is able to locally perceive multimodal sound sources from a scene with visual detection and cross-modality translation. It then contrastively learns a Cross-Modal Sound Source (CMSS) Manifold to semantically disambiguate each source. Finally, we attentively mix their CMSS semantics into a rich audio representation, from which a pretrained audio generator outputs the sound. To model the CMSS manifold, we curate a novel single-sound-source visual-audio dataset VGGS3 from VGGSound. We also design a Sound Source Matching Score to measure localized audio relevance. This is to our knowledge the first work to address V2A generation at the sound-source level. Extensive experiments show that SSV2A surpasses state-of-the-art methods in both generation fidelity and relevance. We further demonstrate SSV2A’s ability to achieve intuitive V2A control by compositing vision, text, and audio conditions. Our SSV2A generation can be tried and heard at https://ssv2a.github.io/SSV2A-demo.
@@ 2.1 Vision-to-Audio Generation
Early V2A methods [41, 6, 68, 18, 5, 7] train a source-specific V2A model on each audio class and cannot generalize to open-domain V2A synthesis. As a precursor, recent SpecVQGAN [22] learns a discrete neural codec [55, 12] of source-agnostic audio features and autoregressively generates audio codes with a Transformer [56]. Following SpecVQGAN, Im2Wav [50] further details its audio codec into low-level and high-level features. Concurrently, MaskVAT [42] leverages a pretrained codec DAC [27] and predicts audio tokens with a Masked Generative Transformer [2]. Another line of methods employ Diffusion [21] models. CLIPSonic-IQ [10] queries CLIP [43] to condition its Diffusion process. Diff-Foley [34] contrastively learns a temporally-aligned visual-audio prior to guide video-audio synchronization. Concurrent method Draw an Audio [63] leverages loudness signal, text caption, and masked video conditions simultaneously. Very recently, some methods bridge visual conditions to the prior of a pretrained audio generator for efficient V2A learning. V2A-Mapper [57] maps CLIP embeddings to CLAP [11] space, from which a pretrained AudioLDM [31] model synthesizes the audio signal. Seeing and Hearing [62] aligns ImageBind [15] visual embeddings to AudioLDM. The concurrent FoleyCrafter [64] devises a timestamp predictor to enhance synchronization during bridging. Very recently, FRIEREN [58] explores V2A generation with Rectified Flow Matching [33]. Existing methods only condition on global visual scenes for V2A synthesis. To the best of our knowledge, a sound source-aware V2A method that accepts multimodal conditions does not exist yet.
@@ 3.2 Cross-Modal Sound Source Manifold
To avoid the loss of rich semantics from CLIP and CLAP due to small training data, we employ a Cross-Modal Contrastive Mask Regularization (CCMR) mechanism to regulate the contrastive guidance $\mathcal{L}_{c}$ defined in Eq. 3. For each batch, we compute a CLIP-CLIP similarity matrix $\mathbf{M}_{\text{CLIP}}$ and a CLAP-CLAP similarity matrix $\mathbf{M}_{\text{CLAP}}$ per entry as:
@@
The CCMR mask sees that high-similarity visual-audio pairs agreed by both CLIP and CLAP are suppressed from participating in the contrastive guidance $\mathcal{L}_{c}$, even when the visual and audio CMSS embeddings are from different sound sources. As a result, these pairs are no longer subject to a high contrastive penalty, and the original semantics from CLIP/CLAP are kept between these samples despite the lack of a large training set.
@@ 3.3 Sound Source Remixer
We model $\psi\left(\cdot\right)$ with a stack of self-attention layers and learn it from visual-audio pairs in VGGSound. The visual sources are perceived from each video’s central frame following Sec. 3.1. Each token sequence $\left\{\mathbf{x}_{m}\right\}$ is zero padded to a fixed length of $M=64$. To enhance generation diversity, a Classifier-free Guidance [20] is applied during training by randomly zeroing out tokens. We replace the classic attention with Efficient Attention [51] and detail this architecture in Sec. A.2.2 During inference, we set $\mathbf{v}=\mathbf{0}$ for sound source conditions from audio modality.
@@
Recall in Sec. 3.2 that we can also obtain a CLAP embedding $\mathbf{a}_{\text{src}}=\chi\left(\mathbf{e}\right)$ for each sound source through the manifold’s reconstructor module. $\mathbf{a}_{\text{src}}$ can be regarded as the source-wise audio semantics generated by our method, i.e. a set of “track semantics”. As one of our objectives for $\mathbf{a}_{\text{mix}}$ is to have high relevance to each sound source, $\left\{\mathbf{a}^{m}_{\text{src}}\right\}$ are recycled to iteratively guide the generation of $\mathbf{a}_{\text{mix}}$. This mechanism, termed Cycle Mix, is illustrated in Fig. 3 (c) and described as Algorithm 1.
@@
So far, the Sound Source Remixer learns an image-to-audio task. We adapt it to the video-to-audio task with a downstream Temporal Aggregation (TA) function $\omega\left(\cdot\right)$ depicted in Fig. 3 (d). We evenly extract 64 frames along time from one video and generate a CLAP embedding for each of them. Each embedding is then positionally embedded with its timestamp. $\omega\left(\cdot\right)$ learns to fuse these embeddings into a temporally-aligned CLAP audio representation $\mathbf{a}$ with the following loss:
@@ 4.1 Experimental Setup
We measure generation quality objectively from two perspectives: fidelity and relevance. For generation fidelity, we adopt the Fréchet Audio Distance (FAD) [45] with an open-source implementation [53] to obtain two metrics, V-FAD and C-FAD, respectively from VGGish [45] and CLAP [11] models. FAD measures the closeness of ground-truth and generated audio feature distributions with a pretrained extractor. A low FAD score reflects high generation fidelity. For generation relevance, we first adopt the CLIP-Score (CS) which maps an audio’s CLAP embedding to the CLIP image space with a pretrained Wav2CLIP [60] model to compare its similarity with the paired image. For multi-source image-audio pairs in VGG-SS, we compute the averaged CS between each sound source image and the paired audio. We compute CS on global images in other tests. A high CS score represents high generation relevance.
@@
We observe that the CS relevance comparison, by mapping audio features to image domain, causes loss of audio information. As a result, our method often outperforms Oracle AudioLDM generations in CS scoring from Tab. 1. We propose a novel metric, Sound Source Matching Score (SSMS), that operates directly in the audio domain for clearer relevance evaluations. SSMS adopts an audio classifier, BEATs [8], to respectively predict $N$ localized sound source labels for ground-truth and generated audios. We regard intersected labels from the predicted sets as true positives, the difference of ground-truth against generation as false negatives and the difference of generation against ground-truth as false positives. SSMS is computed as the F1 score of these statistics. We set $N=10$ throughout experiments and show in Tab. 1 that SSMS distinguishes generation relevance more clearly than CS.
@@
To further establish the soundness of our method, we conduct a subjective listening test with 20 human evaluators. We randomly sample 40 central video frames from AudioSet Strong [19] and AVSBench [67], generating 10-second audio clips with each method. The test participants are asked to rate 20 of them for fidelity without visual cues. They then rate 20 samples for relevance given the visual conditions. We collect the ratings on a 5-point scale and compute the Mean Opinion Score (MOS) [48] to measure generation fidelity and relevance. Please see Appendix A.5 for the evaluation setup.
@@ 4.2 Baseline Evaluations
We compare our generator with four V2A methods: V2A-Mapper [57], Diff-Foley [34], Seeing and Hearing (S&H) [62] and Im2Wav [50]. Currently, V2A-Mapper holds state-of-the-art generation fidelity and relevance. These methods require different visual conditions. For fairness, we modify some methods following Appendix A.1 but still keep their original versions in Tab. 1.
Objective Results.
As illustrated in Tab. 1, our method achieves superior performance in most objective metrics for both in-distribution and out-of-distribution tests. For single-source generation, our SSV2A outperforms baselines in generation relevance and stays in top 2 for generation fidelity. For multi-source generation, SSV2A is superior in all metrics. Surprisingly, SSV2A achieves a higher CS in generation relevance than the Oracle baseline, which is assumed to have optimal performance for V2A methods involving AudioLDM. This effect is no longer observed in SSMS, demonstrating our new metric’s superiority in comparing audio generation relevance. Even though Diff-Foley and S&H-Text are unfair comparisons, SSV2A still surpasses them in both fidelity and relevance.
Subjective Results.
In Tab. 2, we obtain the subjective results as Mean Opinion Score (MOS) from human evaluations. Our method outperforms baselines significantly in both generation fidelity and relevance. Since the random samples are images, we test on the Diff-Foley-Image instead of Diff-Foley because the latter only accepts video inputs. Moreover, we choose to test S&H-Text instead of S&H to obtain the best generation performance Seeing and Hearing can achieve, even though it sees extra text captions.
@@ 4.3 Ablation Study
We conduct several ablation experiments to consolidate our claims in Sec. 3. We also provide a brief analysis on the learned CMSS manifold space to better understand its features in Appendix A.3.
The Effect of CMSS Manifold.
At the beginning of Sec. 3, we mention that SSV2A could learn to perform the V2A task without CMSS disambiguation. In order to prove the benefits of this disambiguation, we perturb the same Sound Source Remixer model with three different generation conditions: without CLIP embeddings, without CMSS embeddings, and with both embeddings. We train them on the same VGGSound data and evaluate the results with VGG-SS tests in Tab. 3. A significant performance drop is observed in both generation fidelity and relevance when the CMSS conditioning is suppressed. This ablation confirms that CMSS disambiguation benefits our V2A task.
The Effect of CCMR.
Recall that in Eq. 7 we have defined a hyperparameter $\alpha$ to control CCMR’s behavior. When $\alpha=0$, the mask becomes an identity matrix and CCMR is stifled. We train the same CMSS manifold modules under four settings of $\alpha$ and conduct VGG-SS tests. Tab. 4 shows that with CCMR, we can enrich the CMSS semantics to benefit downstream generation. However, setting $\alpha$ to higher values degrades generation quality.
The Effect of Cycle Mix.
The Cycle Mix algorithm has two adjustable parameters: the number of iterations and the the sampling size of remixed CLAP embeddings in each iteration. We conduct 8 VGG-SS tests to observe the effect of Cycle Mix parameters in Sec. A.4.1. The optimal setup is chosen as 64 iterations and 64 samples, which is used throughout other experiments.
@@ 4.4 Multimodal Sound Source Composition
Visual Composition.
SSV2A can generate realistic audio by composing visual sound sources. The result respects the supplied sources to render a convincing audio scene. For instance, we can synthesize a “motorbike riders laughing” audio from pictures of a motorbike and a laughing man.
Visual-Text Composition.
SSV2A can further control the V2A generation with textual semantics. For example, we can supply a “motorbike” image and obtain a seaside riding audio with the text prompt “seaside”.
Visual-Audio Composition.
We can achieve a similar style control with audio semantics. For example, we can accompany a “boat pier” image with a “talking” audio to synthesize audio of a busy pier.
Visual-Text-Audio Composition.
Putting it altogether, we can synthesize audio with all three modalities involved. To test this feature, we have successfully produced a “coastline motorcycle racing” audio with a motorcycle image, a “crowd cheering” text, and a “beach” audio.
@@ 5 Limitations and Conclusion
In this work, we explore the feasibility of learning a sound source-aware V2A generator, SSV2A, that supports multimodal conditioning. By explicitly modeling the source disambiguation process with a contrastive cross-modal manifold on single-source visual-audio pairs, we are able to significantly boost our SSV2A’s generation fidelity and relevance. Consequently, SSV2A achieves state-of-the-art V2A performance in both objective and subjective evaluations. Moreover, we demonstrate the intuitive control of our generator in various composition experiments of sound source conditions from vision, text, and audio. To accompany the learning of our sound source disambiguation, we curate a new single-sound-source visual-audio dataset with 106K high-quality visual-audio pairs, namely VGGS3. Additionally, we contribute a novel Sound Source Matching Score that measures audio generation relevance more clearly than the existing CLIP-Score. Two limitations exist in SSV2A. First, we address the video-to-audio synchronization in SSV2A with a naive Temporal Aggregation module. Existing V2A works [34, 42, 64] show that temporal alignment is a nontrivial problem due to the sparsity of synchronization cues in both time and space [24, 23]. Second, we observe that SSV2A is less sensitive to audio conditioning than visual or text inputs. We suspect that this phenomenon is due to the lack of CLIP semantics when the Sound Source Remixer is prompted with audio conditions. In future works, we aim to further investigate these issues.

### v3

# Gotta Hear Them All: Sound Source Aware Vision to Audio Generation

@@ Abstract
Vision-to-audio (V2A) synthesis has broad applications in multimedia. Recent advancements of V2A methods have made it possible to generate relevant audios from inputs of videos or still images. However, the immersiveness and expressiveness of the generation are limited. One possible problem is that existing methods solely rely on the global scene and overlook details of local sounding objects (i.e., sound sources). To address this issue, we propose a Sound Source-Aware V2A (SSV2A) generator. SSV2A is able to locally perceive multimodal sound sources from a scene with visual detection and cross-modality translation. It then contrastively learns a Cross-Modal Sound Source (CMSS) Manifold to semantically disambiguate each source. Finally, we attentively mix their CMSS semantics into a rich audio representation, from which a pretrained audio generator outputs the sound. To model the CMSS manifold, we curate a novel single-sound-source visual-audio dataset VGGS3 from VGGSound. We also design a Sound Source Matching Score to measure localized audio relevance. By addressing V2A generation at the sound-source level, SSV2A surpasses state-of-the-art methods in both generation fidelity and relevance as evidenced by extensive experiments. We further demonstrate SSV2A’s ability to achieve intuitive V2A control by compositing vision, text, and audio conditions. Our generation can be tried and heard at https://ssv2a.github.io/SSV2A-demo.
@@ 2.1 Vision-to-Audio Generation
Early V2A methods [44, 6, 73, 20, 5, 7] train a source-specific V2A model on each audio class and cannot generalize to open-domain V2A synthesis. As a precursor, recent SpecVQGAN [24] learns a discrete neural codec [58, 14] of source-agnostic audio features and autoregressively generates audio codes with a Transformer [59]. Following SpecVQGAN, Im2Wav [53] further details its audio codec into low-level and high-level features. MaskVAT [45] leverages a pretrained codec DAC [29] and predicts audio tokens with a Masked Generative Transformer [2]. Another line of methods employ Diffusion [23] models. CLIPSonic-IQ [12] queries CLIP [46] to condition its Diffusion process. Diff-Foley [37] contrastively learns a temporally-aligned visual-audio prior to guide video-audio synchronization. Draw-an-Audio [66] leverages loudness signal, text caption, and masked video conditions simultaneously. More recently, some methods bridge visual conditions to the prior of a pretrained audio generator for efficient V2A learning. V2A-Mapper [60] maps CLIP embeddings to CLAP [13] space, from which a pretrained AudioLDM [34] model synthesizes the audio signal. V2A-SceneDetector [67] extends V2A-Mapper to multi-scene video with a detection module. Seeing and Hearing [65] aligns ImageBind [17] visual embeddings to AudioLDM. FoleyCrafter [69] devises a timestamp predictor to enhance synchronization during bridging. Very recently, FRIEREN [61] and MMAudio [9] explore V2A generation with Rectified Flow Matching [36]. Most existing methods condition on global visual scenes for V2A synthesis. Some recent works [32] [31] leverage pixel-level conditions for V2A synthesis, partially describing visual sounding objects. In reality, human perceive object-level sound sources across modalities and time [38]. Such a sound source-aware V2A generator remains uninvestigated.
@@ 3.2 Cross-Modal Sound Source Manifold
To avoid the loss of rich semantics from CLIP and CLAP due to small training data, we employ a Cross-Modal Contrastive Mask Regularization (CCMR) mechanism to weaken the contrastive guidance $\mathcal{L}_{c}$ defined in Eq. 3 for similar cross-pair audio-visual samples. For each batch, we compute a CLIP-CLIP similarity matrix $\mathbf{M}_{\text{CLIP}}$ and a CLAP-CLAP similarity matrix $\mathbf{M}_{\text{CLAP}}$ per entry as:
@@ 3.3 Sound Source Remixer
We model $\psi\left(\cdot\right)$ with a stack of self-attention layers and learn it from visual-audio pairs in VGGSound. The visual sources are perceived from each video’s central frame following Sec. 3.1. Each token sequence $\left\{\mathbf{x}_{m}\right\}$ is zero padded to a fixed length of $M=64$. To enhance generation diversity, a Classifier-free Guidance [22] is applied during training by randomly zeroing out tokens. We replace the classic attention with Efficient Attention [54] and detail this architecture in Sec. A.2.2. During inference, we set $\mathbf{v}=\mathbf{0}$ for sound source conditions from audio modality.
@@
Recall in Sec. 3.2 that we can also obtain a CLAP embedding $\mathbf{a}_{\text{src}}=\chi\left(\mathbf{e}\right)$ for each sound source through the manifold’s reconstructor module. $\mathbf{a}_{\text{src}}$ can be regarded as a set of source-wise audio semantics generated by our method. As one of our objectives for $\mathbf{a}_{\text{mix}}$ is to have high relevance to each sound source, $\left\{\mathbf{a}^{m}_{\text{src}}\right\}$ are recycled to iteratively guide the generation of $\mathbf{a}_{\text{mix}}$. This mechanism, termed Cycle Mix, is illustrated in Fig. 3 (c) and Algorithm 1.
@@
So far, the Sound Source Remixer learns an image-to-audio task. Following V2A-Mapper [60], we adapt it to the video-to-audio task with a downstream Temporal Aggregation (TA) function $\omega\left(\cdot\right)$ depicted in Fig. 3 (d). Instead of averaging the frame-wise semantics, we learn a nonlinear $\omega\left(\cdot\right)$. We evenly extract 64 frames along time from one video and generate a CLAP embedding for each of them. Each embedding is then positionally embedded with its timestamp. $\omega\left(\cdot\right)$ learns to fuse these embeddings into a temporally-aligned CLAP audio representation $\mathbf{a}$ with the following loss:
@@
Method WMAO $\downarrow$
Top-1 Top-5
Diff-Foley 1.227 1.127
FoleyCrafter 1.247 1.167
V2A-Mapper 1.299 1.197
Ours w/o TA 1.292 1.205
Ours 1.243 1.172
Table 3: Synchronization tests on AVSync. Red, orange, and brown mark first, second, and third placements.
@@ 4.1 Experimental Setup
We measure generation quality objectively from two perspectives: fidelity and relevance. For generation fidelity, we adopt the Fréchet Audio Distance (FAD) [48] with an open-source implementation [56] to obtain two metrics, V-FAD and C-FAD, respectively from VGGish [48] and CLAP [13] models. FAD measures the closeness of ground-truth and generated audio feature distributions. A low FAD score reflects high generation fidelity. For generation relevance, we first adopt the CLIP-Score (CS) which maps an audio’s CLAP embedding to the CLIP image space with a pretrained Wav2CLIP [63] model to compare its similarity with the paired image. For multi-source image-audio pairs in VGG-SS, we compute the averaged CS between each sound source image and the paired audio. We compute CS on global images in other tests. A high CS score represents high generation relevance.
@@
We observe that the CS relevance comparison, by mapping audio features to image domain, causes loss of audio information. As a result, our method often outperforms Oracle AudioLDM generations in CS scoring from Tab. 1. We propose a novel metric, Sound Source Matching Score (SSMS), that adopts an audio classifier BEATs [8] to respectively predict $N$ localized sound source labels for ground-truth and generated audios. We regard intersected labels from the predicted sets as true positives, the difference of ground-truth against generation as false negatives and the difference of generation against ground-truth as false positives. SSMS is computed as the F1 score of these statistics. We set $N=10$ throughout experiments and show in Tab. 1 that SSMS distinguishes generation relevance more clearly than CS.
Weighted Mean Absolute Offset.
To prove the efficacy of our TA mechanism for video-audio synchronization, we employ SynchFormer [25] to predict the temporal Weighted Mean Abolute Offset (WMAO) in seconds between the original video and the generated audio on the AVSync15 [68] dataset. We weight and sum the top-k predictions of SyncFormer with their confidence scores. A lower WMAO perceives smaller drifts between video-audio signals, indicating higher synchronization.
@@
Following recent works [60, 69, 45], we conduct a subjective listening test with 20 human evaluators. We randomly sample 40 central video frames from AudioSet Strong [21] and AVSBench [72], generating 10-second audio clips with each method. The test participants are asked to rate 20 of them for fidelity without visual cues. They then rate 20 samples for relevance given the visual conditions. We collect the ratings on a 5-point scale and compute the Mean Opinion Score (MOS) [51] to measure generation fidelity and relevance. Please see Appendix A.5 for the evaluation setup.
@@ 4.2 Baseline Evaluations
We compare our generator with four V2A methods: V2A-Mapper [60], Diff-Foley [37], Seeing and Hearing (S&H) [65] and Im2Wav [53]. Currently, V2A-Mapper holds state-of-the-art generation fidelity and relevance. These methods require different visual conditions. For fairness, we modify some methods following Appendix A.1 but still keep their original versions in Tab. 1. For WMAO tests, we compare our method with video-to-audio methods V2A-Mapper, Diff-Foley and FoleyCrafter [69].
Objective Results. As illustrated in Tab. 1, our method achieves superior performance in most objective metrics for both in-distribution and out-of-distribution tests. For single-source generation, our SSV2A outperforms baselines in generation relevance and stays in top 2 for generation fidelity. For multi-source generation, SSV2A is superior in all metrics. Surprisingly, SSV2A achieves a higher CS in generation relevance than the Oracle baseline, which is assumed to have optimal performance for V2A methods involving AudioLDM. This effect is no longer observed in SSMS, demonstrating our new metric’s superiority in comparing audio generation relevance. Even though Diff-Foley and S&H-Text are unfair comparisons, SSV2A still surpasses them in both fidelity and relevance.
On the other hand, our method generates competitive video-audio synchronization as shown by the WMAO tests in Tab. 3. The ablation in TA mechanism shows that it is key to SSV2A’s temporal alignment capability. Moreover, the results indicate that our nonlinear TA function performs better than V2A-Mapper’s linear setup.
Subjective Results. In Tab. 2, we obtain the subjective results as Mean Opinion Score (MOS) from human evaluations. Our method outperforms baselines significantly in both generation fidelity and relevance. Since the random samples are images, we test on the Diff-Foley-Image instead of Diff-Foley because the latter only accepts video inputs. Moreover, we choose to test S&H-Text instead of S&H to obtain the best generation performance Seeing and Hearing can achieve, even though it sees extra text captions.
@@ 4.3 Ablation Study
We conduct several ablation experiments to consolidate our claims in Sec. 3. We also provide an analysis on the learned CMSS manifold space and more ablations in Appendices A.3 and A.4.
Effect of CMSS Manifold. At the beginning of Sec. 3, we mention that SSV2A could learn to perform the V2A task without CMSS disambiguation. In order to prove the benefits of this disambiguation, we perturb the same Sound Source Remixer model with three different generation conditions: without CLIP embeddings, without CMSS embeddings, and with both embeddings. We train them on the same VGGSound data and evaluate the results with VGG-SS tests in Tab. 4. A significant performance drop is observed in both generation fidelity and relevance when the CMSS conditioning is suppressed. This ablation confirms that CMSS disambiguation benefits our V2A task.
Effect of CCMR. Recall that in Eq. 7 we have defined a hyperparameter $\alpha$ to control CCMR’s behavior. When $\alpha=0$, the mask becomes an identity matrix and CCMR is stifled. We train the same CMSS manifold modules under four settings of $\alpha$ and conduct VGG-SS tests. Tab. 5 shows that with CCMR, we can enrich the CMSS semantics to benefit downstream generation. However, setting $\alpha$ to higher values degrades generation quality.
@@ 4.4 Multimodal Sound Source Composition
Visual Composition. SSV2A can generate realistic audio by composing visual sound sources. The result respects the supplied sources to render a convincing audio scene. For instance, we can synthesize a “motorbike riders laughing” audio from pictures of a motorbike and a laughing man.
Visual-Text Composition. SSV2A can further control the V2A generation with textual semantics. For example, we can supply a “motorbike” image and obtain a seaside riding audio with the text prompt “seaside”.
Visual-Audio Composition. We can achieve a similar style control with audio semantics. For example, we can accompany a “boat pier” image with a “talking” audio to synthesize audio of a busy pier.
Visual-Text-Audio Composition. Putting it altogether, we can synthesize audio with all three modalities involved. To test this feature, we have successfully produced a “coastline motorcycle racing” audio with a motorcycle image, a “crowd cheering” text, and a “beach” audio.
Comparison with Cascaded Composition. One could cascade video-to-audio and text-to-audio models to generate source-specific audio clips and overlay them together. We show that our SSV2A composition offers crucial interaction, context and style awareness when integrating sound sources, which are unachievable with this cascaded approach. Please see Appendix A.6 for details.
@@ 5 Limitations and Conclusion
In this work, we explore the feasibility of learning a sound source-aware V2A generator, SSV2A, that supports multimodal conditioning. By explicitly modeling the source disambiguation process with a contrastive cross-modal manifold on single-source visual-audio pairs, we are able to significantly boost our method’s generation fidelity and relevance. Consequently, SSV2A achieves state-of-the-art V2A performance in both objective and subjective evaluations. Moreover, we demonstrate the intuitive control of our generator in various composition experiments of sound source conditions from vision, text, and audio. To accompany the learning of our sound source disambiguation, we curate a new single-sound-source visual-audio dataset VGGS3. Additionally, we contribute a novel Sound Source Matching Score that measures audio generation relevance more clearly than the existing CLIP-Score. Two limitations exist in SSV2A. First, we address the video-to-audio synchronization in SSV2A with a naive temporal module. Existing works [26, 25] show that temporal alignment is a nontrivial problem due to the sparsity of synchronization cues in both time and space. Second, we observe that SSV2A is less sensitive to audio conditioning than visual or text inputs. We suspect that this phenomenon is due to the lack of CLIP semantics when the Sound Source Remixer is prompted with audio conditions. Please see Appendix A.7 for our proposed solutions for these limitations.